OpenStudy (anonymous):

Does anyone have Algebra 2? Solve the system by substitution. 2x+y=-11 3x-4y=11

OpenStudy (dmndlife24):

give me one sec

OpenStudy (anonymous):

Okay, thank you!

rishavraj (rishavraj):

multiply equation 1 from 3 and equation 2 from w.. and substract the two....

OpenStudy (anonymous):

are you trying to find X and Y?

rishavraj (rishavraj):

its 2 instead of w

OpenStudy (anonymous):

Yes, x and y

OpenStudy (anonymous):

what? @rishavraj that makes no sense.

OpenStudy (dmndlife24):

x = -3

OpenStudy (anonymous):

lets call the formulas A ans B ok

OpenStudy (anonymous):

I was confused with what he said too. I keep getting a decimal so I know what I'm doing is wrong. And @dmndlife24 then y would be -5 correct?

OpenStudy (dmndlife24):

y = 5

OpenStudy (dmndlife24):

wait y = -5

OpenStudy (anonymous):

we'll start on A. first you need to get Y alone. to do this subtract 2X y=-11-2X now we can place this in formula B in place for the Y

OpenStudy (dmndlife24):

just try plugging it in back into the original equation

rishavraj (rishavraj):

see use elimination method 1st equation is 2x + y = -11 .. multiply it with 3.. so it is 6x + 3y = -33 ....... (i) 2nd equation 3x - 4y = 11 ...multiply it with 2 will have 6x - 8y = 22 ...... (ii) now just substract the two new equations ...x would be eliminated....

OpenStudy (anonymous):

Okay, I figured out where I messed up now! Thank you so much And no @rishavraj I was supposed to use the substitution method.

OpenStudy (anonymous):

too many people.

rishavraj (rishavraj):

@cheymariee i am sorry I didn't see tht...sorry